# Visual Scratchpads: Enabling Global Reasoning in Vision

### v1

@@ Abstract
Modern vision models have achieved remarkable success in benchmarks where local features provide critical information about the target. There is now a growing interest in solving tasks that require more global reasoning, where local features offer no significant information. These tasks are reminiscent of the connectivity tasks discussed by Minsky and Papert in 1969, which exposed the limitations of the perceptron model and contributed to the first AI winter. In this paper, we revisit such tasks by introducing four global visual benchmarks involving path findings and mazes. We show that: (1) although today’s large vision models largely surpass the expressivity limitations of the early models, they still struggle with the learning efficiency; we put forward the ‘globality degree’ notion to understand this limitation; (2) we then demonstrate that the picture changes and global reasoning becomes feasible with the introduction of ‘visual scratchpads’; similarly to the text scratchpads and chain-of-thoughts used in language models, visual scratchpads help break down global tasks into simpler ones; (3) we finally show that some scratchpads are better than others, in particular, ‘inductive scratchpads’ that take steps relying on less information afford better out-of-distribution generalization and succeed for smaller model sizes.
@@ 1 Introduction
Modern computer vision models, as well as text models, are often pre-trained on vast datasets encompassing much of the knowledge available on the internet. While this has led to impressive capabilities, there is growing concern that these models may make decisions based on shallow, local information rather than engaging in deep, complex reasoning. Evidence suggests that many of these models function primarily through retrieval, acting as blurry, compressed versions of the Internet. These models excel at smooth interpolation within this encoded knowledge but often fail to grasp the underlying logic and complexity of the real world. Unfortunately, the community lacks benchmarks that rigorously test a model’s ability to perform global reasoning and multi-step problem-solving in the visual domain. Instead, most common visual benchmarks are limited to tasks that can be solved with superficial cues and local features. In this work, we aim to close this gap by exploring whether current models are capable of learning tasks that require deep, multi-step global processing.
@@
In order to address this, it is crucial to define the characteristics of global visual tasks. In contrast to local tasks, where a small subset of pixels—typically organized into patches—is sufficient to achieve better-than-random accuracy, global tasks require a more holistic understanding of the entire visual scene. For example, in ImageNet classification (Deng et al., 2009), a single patch containing cat whiskers significantly increases the likelihood that the model will classify the image as a cat. This reliance on local features is further exemplified by the effectiveness of drastic image cropping in object-centric datasets, where self-supervised models such as DINO (Caron et al., 2021) employ aggressive multi-crop strategies, sometimes cropping as much as 90% of the image which empirically improves the performance. Humans, in contrast, do not rely solely on local information; for instance, when driving a car, it is insufficient to focus only on the view directly in front of the vehicle. A competent driver must recognize multiple visual objects in the environment and consider their complex behaviors before making decisions. Yet, using such complex real-world tasks, like autonomous driving, to study model learning is impractical due to their complexity and unpredictability. Instead, we need interpretable and deterministic tasks with well-defined data generation processes to assess the reasoning ability of the models.
Figure 1: The cover of the 2017 edition of Perceptrons by Minsky & Papert (1969).
@@
To address this need for visual tasks with global multi-step reasoning, we propose four simple datasets, some reminiscent of the connectivity task that played a significant role in the AI winter. The connectivity task was studied in the book Perceptrons (Minsky & Papert, 1969), whose cover also illustrated a visual version of the task. Figure 1 shows the 2017 edition of the book, which closely resembles the cover of the 1972 edition. Minsky & Papert (1969) proved that single-layer perceptrons cannot distinguish the two figures based on connectivity due to expressivity issues. Our tasks are closely related to these early studies as they require an understanding of the connectivity concept. In particular, we propose a graph connectivity task, the task of determining the number of strings in an image, and a maze solvability task where we consider rectangular and circular mazes. We argue that these tasks possess the key ingredients for testing global reasoning. They are inherently global because understanding a small portion of the graph or maze offers no meaningful insight into the final label (whether the structure is connected or not). At the same time, their data generation processes are fully controllable and deterministic, allowing for straightforward manipulation of task complexity by simply increasing the number of nodes in the graph or cells in the maze. We have also developed more visually engaging variants, namely the strings task and the circular mazes to enhance the visual aspect. These datasets enable us to simultaneously test both reasoning and visual recognition abilities, which is the core objective of this paper.
Despite the increased expressivity of modern vision models compared to the perceptrons discussed by Minsky & Papert (1969), current models still struggle with global tasks. While they can solve simple and small graph connectivity and maze problems, their performance rapidly declines to the accuracy of a random model as the tasks become more complex. This deterioration occurs regardless of the model’s size, the task used, or whether a pre-trained checkpoint is employed (see Section 4.1). To remedy this issue, we put forward the notion of visual scratchpads. Similar to the scratchpad idea used in the text domain (Nye et al., 2021), a visual scratchpad is a single frame or a sequence of frames that depict the underlying reasoning behind the label of a sample, e.g., the existence of a path between the source and sink nodes in a maze. The model is supervised with the scratchpads during training and has to generate them at test time. The scratchpad acts as a guide, showing the model how to decompose the global problem into simpler subproblems, such as coloring two nodes at a time in a graph or a few cells at a time in a maze. Interestingly, we find that even using a one-shot single-frame scratchpad, which only provides a visualization of the final solution boosts performance significantly making the model capable of learning most of the considered tasks. Furthermore, the model exhibits a hierarchical “staircase” learning behavior, learning the solution incrementally during training, even though it was trained to generate only the final solution in a single shot. Moreover, for the multi-frame scratchpad, we propose a model that generates scratchpad frames in an autoregressive, recurrent, and Markovian manner called the inductive scratchpad model. We show that this model outperforms the single-scratchpad model both in in-distribution and out-of-distribution settings thanks to its Markovian modeling and adaptive compute time at inference.
Our work differs from previous efforts in textual scratchpads due to the unique characteristics of visual data. Unlike language, which consists of discrete tokens, vision deals with continuous inputs. Additionally, vision operates within a 2D spatial neighborhood of pixels or patches, in contrast to the linear, 1D structure of text. This difference influences how models generalize to OOD samples (such as generalization to more challenging samples that require more reasoning steps), as more objects can be represented within the same pixel space without requiring additional positional embeddings. These distinct features make vision a particularly interesting and fertile field for applying and extending the scratchpad concept. Here is a summary of this paper’s contributions:
  • Study of locality and globality in computer vision: we analyze the concept of locality/globality in vision, to distinguish between global and local tasks while paying particular attention to the vision Transformers (ViTs, Dosovitskiy et al., 2020), the currently dominant models.
  • Development of datasets inspired by Minsky & Papert’s connectivity task: we revisit the foundational work of Minsky & Papert (1969) by proposing four tasks related to the connectivity concept that require global reasoning and are hard to learn for ViT models of different sizes.
  • Introduction of visual scratchpads for global reasoning: we introduce the visual scratchpad to enable multi-step reasoning in vision models. More specifically,
    • we show that a single-frame scratchpad model can learn visual tasks that were not learnable with the no-scratchpad models irrespective of size and pre-training;
    • we introduce a recurrent model for generating multi-frame scratchpads, namely, the inductive scratchpad model that allows for better in-distribution and out-of-distribution (OOD) generalization thanks to its Markovian modeling and adaptive compute time at inference.
2 Global visual reasoning datasets
Vision models have shown remarkable performance on different tasks including image classification, image segmentation, object detection, etc. However, these mainstream visual tasks have two characteristics in common:
  1. 1.
    Local features in the image are informative. For example, if we consider an image partitioned into a set of patches, there is usually a small subset of patches that provides significant information on the target (e.g., the label).
  2. 2.
    These tasks can be instinctively and instantaneously solved by humans. That is humans do not need to ponder for longer periods of time to solve these tasks. Considering the System 1 / System 2 terminology (Kahneman, 2011), these visual tasks are dealt with by our System 1. In general little or no multi-step chain of entailments is necessary to solve these tasks (e.g., no search).
Despite being common, not all visual tasks share these characteristics. As an example, consider solving a maze, i.e., answering whether two points are connected in a maze or not. Assuming the size of the maze is large enough, humans need some deliberation before solving the maze. Normally, humans would follow paths with a pen on the maze to see where the starting point leads to. Importantly, apart from trivial edge cases where the start and end locations are close, local features are not informative on the maze task. For instance, if only three patches of a maze are given, one cannot solve the maze (determine whether there is a connection) with high probability. Motivated by the latter, we propose the following visual datasets in this paper:
  • Connectivity datasets. Inspired by Minsky & Papert (1969), we consider two datasets based on the notion of connectivity.
    • Cycles task. In this task, $2n$ nodes are drawn randomly (on an invisible circle) in the image. There are also $2n$ edges between these nodes that form either one cycle of size $2n$ or two cycles of size $n$. The task is to determine whether the graph is connected (one cycle, label 1) or not (two cycles, label 0) given an input image. See Figure 2 for an example. In this task, one has reason on at least $n$ nodes and the connections between them to determine the label correctly as any $n-1$ nodes do not provide any information on whether there are two cycles or one. Thus, one can simply increase the complexity of this task by increasing $n$.
    • Strings task. In order to further increase the visual complexity, we consider a dataset consisting of random strings. In each sample, there are either two closed strings or one longer closed string. The dataset generation process for these curves is similar to the cycles task above with the difference that in the strings we do not make the (anchor) nodes visible and also connect them using 3rd-degree Bézier curves which produces continuous strings (see Figure 2). Similar to the cycles task, one can increase the complexity of this task by increasing the number of invisible anchor points $2n$ which leads to longer more entangled strings.
  • Maze solvability. We also consider a maze task in which there are always two connected components, and we have a start/source point (shown in blue) and an end/sink point (shown in red). The source and sink are in the same connected component or not equiprobably. The task is to determine whether they are connected (label 1) or not (label 0). We provide this dataset in a rectangular and a circular version to increase the visual complexity. Examples can be seen in Figure 2. To adjust the complexity of maze datasets, one can modify the size of the maze and hence the number of cells, size of the components, and distance between the source and sink (if connected).
For each task, there exists a natural visual scratchpad that uncovers the underlying reasoning behind the label. For the maze, similar to what humans do we can start coloring from the source cell (e.g., the cell in blue) to see which areas are reachable until reaching the sink cell or the end of the maze region similar to doing a breadth-first search (BFS). This coloring is similar to what humans would naturally do by following the paths from the beginning to see which one (if any) leads to the sink cell. For the cycles task, we can use a similar idea, we can start by coloring one node, and then coloring all of the nodes that are connected to this node (which is either half of the graph or all of the graph). Analogously, for the strings task the visual scratchpad would be coloring one of the strings if there are two strings or coloring the whole string if there is only one. To disambiguate which cycle/string to color, we always color the cycle/string that passes through the rightmost (anchor) node.
Note that the scratchpad can have a single frame format in which the full scratchpad (all coloring done) is shown. The scratchpad can also be generated in multiple frames, i.e., consecutive frames that lead to the final frame. This scratchpad is again analogous to what humans do: starting from one point and coloring progressively. For example, this could be coloring a distance of $10$ when doing the search for the maze problems, and (up to) two (anchor) nodes for the cycles and strings tasks. An example of doing so for cycles task is depicted in Figure 3. We start the coloring from the source node in mazes and from the rightmost (anchor) node in cycles and strings examples. See Appendix D.6 for scratchpad figures for other datasets.
@@ 2 Global visual reasoning datasets
Figure 2: Examples of different tasks. The first row shows the inputs. The second row shows the complete scratchpad (e.g., the target frame in the single-frame scratchpad model and the final frame in the multi-frame scratchpad). In the examples of cycles and strings, we have two different cycles and strings. In the example of the circular maze, the two cells are not connected while in the example of the rectangular maze, they are. In each of the scratchpads, we start coloring from some point until the whole connected part is colored (or the sink cell is reached in the maze examples).
@@
Figure 3: Example of the cycles task showing how the scratchpad can contain several frames. The input image is presented on the left side and then different frames of the visual scratchpad are depicted from left to right ending with the complete image.
2.1 Globality, locality and scratchpads in the visual domain
Here, we further discuss the meaning of locality and globality in the visual domain and the connection of the visual scratchpad to the concept of scratchpad in the text domain (Nye et al., 2021).
@@ 2.1 Globality, locality and scratchpads in the visual domain
Recently, Abbe et al. (2024) proposed the notion of globality degree to explain why some tasks are hard to learn for Transformers and also to explain the effectiveness of scratchpads in the textual domain. Considering input tokens $X_{1},\ldots,X_{n}$ and output $Y$, the globality degree of a task is defined as the minimum number of tokens $k$ such that there exist $k$ tokens $X_{i_{1}},\ldots,X_{i_{k}}$ that along with the histogram of tokens $\hat{P}_{X}$ [^1] In the textual domain, histogram simply refers to reporting how many times each token is appearing regardless of its position (similar to the bag of words). provide significant information on the target $Y$, i.e., $I(X_{i_{1}},\ldots,X_{i_{k}},\hat{P}_{X};Y)=n^{-O_{n}(1)}$ where $I$ is the mutual information. It is further conjectured, with empirical support, that the learning complexity of tasks increases with their globality degree, and Transformers can only learn tasks with a constant globality degree (in $n$) efficiently (polysize model and polymany iterations). We extend this definition to vision tasks learnable by ViTs.
@@
Globality degree (in vision with significant patch regime). Assume images are partitioned into patches $X_{1},\ldots,X_{n}$. We define the globality degree with threshold $\alpha$ of a task as the minimum number $k$ such that there exist patches $X_{i_{1}},\ldots,X_{i_{k}}$ that satisfy $I(X_{i_{1}},\ldots,X_{i_{k}};Y)\geq\alpha$ where $I$ is the mutual information.
In words, this is the least number of patches $k^{*}$ required to obtain an $\alpha$-mutual information with the target. The higher $\alpha$, the more informative these patches are on the target, and the lower $k^{*}$ (for the same $\alpha$), the less global the task is. One may now try to use this measure to characterize which targets are learnable in polynomial time in the number of patches $n$. This requires a closer investigation of the scale of the parameters. To define our asymptotic quantities properly, we assume that the number of patches $n$ is scaling (e.g., as the size of the maze increases, we need a higher resolution image to solve it). [^2] The number of patches in ViT models is usually constant as the images are usually resized unless the image is so fine-grained that a higher resolution is required, e.g., when the number of graph nodes diverges. In this case, the requirement would be to have $\alpha=n^{-O_{n}(1)}$ and $k^{*}=O_{n}(1)$ in order for the whole [^3] I.e., assuming networks with polynomial number of edges trained with a polynomial number of samples/epochs, the overall complexity would be polynomial learning complexity to be polynomial in $n$, i.e., we expect the complexity of learning in this regime to depend polynomially on both $n$ and $1/\alpha$ with an exponent given by $k^{*}$, i.e., $poly(\frac{1}{\alpha},n)^{k^{*}}$ where $k^{*}$ is the globality degree. Our main regime of interest, the regime with significant patches, assumes that patches are large enough, e.g., $P\times P$ sized-patch with $P=\sqrt{n}$, such that only a unique ordering of the patches is valid (if one permutes the patches the new sample does not belong to the distribution’s support w.h.p. as in many computer vision tasks of interest). In this case, the histogram part should not be inserted as done in our current definition. This is why we call this the globality degree in the ‘significant patch regime’, which we consider to to be the right regime to better understand the targets of interest. In the small-patch regime where the patches have a constant size, i.e., $P\times P$ sized-patch with $P=O_{n}(1)$, one would also need to update slightly the definition of the globality degree in order to capture the fact that targets that are permutation invariant may be more easily learnable by the Transformer after dropping the positional embeddings, which results in adding the histogram of the patches to the mutual information as was done in the globality degree for text by Abbe et al. (2024).
@@
Note that a task being “local” by affording a low $k^{*}$ for a significant $\alpha$ does not mean that it does not depend on all the patches, but that these few $k^{*}$ patches are sufficient in order to obtain non-trivial information about the target, and this gives the starting point to learning with a significant edge.
@@
According to the definition above, classical vision tasks such as image classification are “local” as a few patches often provide significant information on the class (e.g., having a patch containing a dog’s ear). Whereas, in our proposed datasets, seeing a few patches often provides no information on the label. Hence, our proposed datasets have a high globality degree, or in short, are rather “global”. For example, in the maze examples, seeing just a few patches from the maze does not help with determining the label. Similarly, for the cycles task of size $2n$, if one only sees the connection between $n-1$ nodes, one cannot have any information on the label.
@@
In order to further support our claim that mainstream vision tasks are local while our proposed tasks are not, we conduct the following experiment. For each sample in a given dataset, we mask the patches with probability $p$ at both training and inference and see the performance of the model with different values of $p$. We perform this experiment for the cycles 12 task and ImageNet. Since the the cycles 12 task is not learnable from scratch we start from a CLIP (Radford et al., 2021) pre-trained ViT-L/14 checkpoint (Fang et al., 2023) for both. The results are shown in Figure 5 (left)  [^4] We use min-max normalized accuracy in the plot, including the random baseline for normalization.. We observe that the model demonstrates good performance on the ImageNet dataset even when 90% of the image is masked while it cannot learn the cycles 12 task once 30% (or more) of the image is masked. The latter shows that the cycles 12 task is a high globality dataset where one needs on average at least 70% of the patches to gain minimal information on the label while ImageNet is a local task where weak learning is possible with only 10% of the patches.
Interestingly, we note that a model trained from scratch was not able to learn the cycles task even when no patch was masked. To better explain this phenomenon, in Figure 5 (right), we compare the performance when initializing with a pre-trained model (Fang et al., 2023) versus a model trained from scratch on the cycles task of varying size. This shows that as the task complexity increases with the number of nodes, none of the models are able to learn the cycles task, meaning that even strong internet-sourced priors in the pre-trained model are not helpful.
@@
Figure 4: Experimental evidence that cycles 12 is a more global task w.r.t. common computer vision benchmarks (e.g. ImageNet). Cycles 12 quickly becomes not learnable when more patches are masked, while ImageNet is still far from random accuracy.

### v2

# Chain-of-Sketch: Enabling Global Visual Reasoning

@@ Abstract
Modern vision models have achieved remarkable success in benchmarks where local features provide critical information about the target. There is now a growing interest in tackling tasks requiring more global reasoning, where local features do not provide significant information. Minsky and Papert put forward such tasks in 1969 with their connectivity study, exposing the limitations of the perceptron model. In this paper, we introduce an expanded set of global visual datasets involving graphs, strings, mazes, and image grids. We show that large vision models still struggle to learn these tasks efficiently. Similarly, state-of-the-art multi-modal LLMs perform poorly on these datasets. We explain this learning inefficiency by means of the ‘globality degree’ measure. To mitigate this, we propose a method called chain-of-sketch (CoS). Similar to the chain-of-thought and scratchpad techniques used in language models, CoS breaks the original task into intermediate visual steps to help learn a complex task. In addition, we show that not all CoS strategies perform equally well. Our key insight is to impose a Markovian structure on the CoS frames. This leads to the introduction of ‘inductive CoS’ which achieves better out-of-distribution generalization and performs well even with smaller models compared to non-inductive variants. [^1] An earlier version appeared on arXiv under the title “Visual Scratchpads: Enabling Global Reasoning in Vision”.
@@ 1 Introduction
Modern computer vision models, as well as text models, are often pre-trained on vast datasets encompassing much of the knowledge available on the internet. While this has led to impressive capabilities, measuring the extent to which these models perform reasoning is still under investigation. Evidence suggests that many of these models, acting as blurry, compressed versions of the Internet, excel at smooth interpolation within their encoded knowledge, but often struggle to grasp underlying logic and extrapolate robustly. Unfortunately, classical visual benchmarks are limited to tasks that can often be tackled with superficial cues and local features. Despite progress in visual and multi-modal reasoning benchmarks (Yue et al., 2023, 2024; Hao et al., 2025), there is a need for datasets that rigorously test global reasoning and multi-step visual problem-solving. In this work, we aim to bridge this gap by exploring when and how models are capable of learning tasks that require multi-step global processing of the input.
@@
To that end, it is crucial to define the characteristics of global visual tasks. In contrast to local tasks, where a small subset of pixels—typically organized into patches—is sufficient to achieve better-than-random accuracy, global tasks require a more holistic understanding of the entire visual scene. For example, in ImageNet classification (Deng et al., 2009), a single patch containing cat whiskers significantly increases the likelihood that the model will classify the image as a cat. This reliance on local features is further exemplified by the effectiveness of drastic image cropping in object-centric datasets, where self-supervised models such as DINO (Caron et al., 2021) employ aggressive multi-crop strategies, sometimes cropping as much as 90% of the image, which empirically improves performance. Humans, in contrast, do not rely solely on local information; for instance, when driving a car, it is insufficient to focus only on the view directly in front of the vehicle. A competent driver must recognize multiple visual objects in the environment and consider their complex behaviors before making decisions. Yet, using such complex real-world tasks, like autonomous driving, to study learning is impractical due to their complexity and unpredictability. Instead, we need interpretable and deterministic tasks with well-defined data generation processes to assess the reasoning ability of the models.
Figure 1: The cover of the 2017 edition of Perceptrons by Minsky and Papert (1969).
@@
To address this need for visual tasks with global multi-step reasoning, we propose five datasets, some reminiscent of the connectivity task of Minsky and Papert (1969) that played a significant role in the AI winter (see Figure 1). In particular, we propose two connectivity-based tasks (cycles and strings) and two maze solvability tasks (rectangular and circular mazes). In addition, we re-purpose the pointer value retrieval (PVR) task (Zhang et al., 2021b; Abbe et al., 2022a) in a visual setting. These tasks are inherently global because understanding a small portion of the input offers no meaningful information into the final label (e.g., whether the structure is connected or not). Further, their data generation processes are fully controllable and deterministic, allowing for straightforward manipulation of the task’s complexity (e.g., by changing the size of the maze). These datasets enable us to simultaneously assess both reasoning and visual recognition abilities, which is the core objective of this paper. Interestingly, we find that even state-of-the-art multi-modal LLMs such as GPT-4o (Hurst et al., 2024), o3, and o4-mini (OpenAI, 2025) struggle on these datasets (see Section 2.2), barely performing better than random, which demonstrates the need for such global visual reasoning datasets.
1.1 Contributions
  • Exploration of locality and globality in the visual domain: We analyze the concept of locality/globality in vision, providing a measure of the task’s globality degree to better understand the learning complexity of models such as Vision Transformers (ViTs) (Dosovitskiy et al., 2020).
  • Datasets to investigate the limitations of global visual reasoning: We propose a set of tasks involving graphs, strings, and mazes based on the connectivity idea and a pointer value retrieval task. These tasks are special as (1) no small subset of patches is sufficient to even weakly learn the target (and a more global multi-step reasoning is needed); (2) the complexity of the tasks can be scaled with task hyperparameters; (3) ViT models struggle to learn these tasks as the task complexity grows. We provide an explanation of these limitations using the globality degree measure (Abbe et al., 2024).
  • Chain-of-sketch for global reasoning: Similar to the chain-of-thought in text (Wei et al., 2023), we develop chain-of-sketch (CoS) techniques to enable multi-step reasoning in vision models. CoS refers to the idea of learning intermediate visual targets that facilitate learning the original target, especially for global targets. Specifically, (1) we show that a single-frame CoS model can learn visual tasks that were not learnable without CoS; (2) we introduce a recurrent model for generating multi-frame CoS, the inductive CoS, that uses Markovian modeling of the intermediate steps and adaptive compute time at inference. We show that inductive CoS affords stronger reasoning capabilities on the above tasks by improving the OOD generalization (length generalization) and by enabling smaller models to learn despite their failure with the non-inductive CoS.
@@ 1.1 Contributions
(e) Image-PVR
Figure 2: Examples of different tasks. The first row shows the inputs; the second row shows the complete sketch (i.e., the target frame in single-frame CoS or the final frame in multi-frame CoS). In the graph and string tasks, there are one or two connected components (two shown). In the maze tasks, there are always two connected components, but the start and end nodes may or may not be connected (both cases shown). In the PVR task, the label is the parity of the airplane class in the indicated row (00 in this example).
2 Global visual reasoning datasets
Vision models have shown remarkable performance on a range of tasks, including image classification, image segmentation, and object detection. However, mainstream vision datasets, including reasoning benchmarks described in Section 5.1, have two characteristics in common:
  1. 1.
    Local features in the image are informative. For example, if we consider an image partitioned into a set of patches, there is usually a small subset of patches that provides significant information about the target (e.g., the label).
  2. 2.
    These tasks can be instinctively and instantaneously solved by humans. That is, humans do not need to ponder for long periods of time to solve these tasks. Considering the System 1 / System 2 terminology (Kahneman, 2011), these visual tasks are handled by our System 1. In general, little or no multi-step chain of entailments is necessary to solve these tasks.
However, not all visual tasks share these characteristics. For instance, consider solving a maze, i.e., answering whether two points in a maze are connected or not. Assuming the size of the maze is large enough, humans require some reflection before solving the maze. Normally, humans would trace the paths on the maze with a pen to see where the starting point leads. Importantly, apart from trivial edge cases where the start and end locations are close, local features are not informative for the maze task. For example, if only a few patches of a maze are given, one cannot solve it with high probability. Motivated by the latter, we propose the following visual datasets in this paper:
  • Connectivity datasets. Inspired by Minsky and Papert (1969), we consider two datasets based on the notion of connectivity.
    • Cycles task. In this task, $2n$ nodes are drawn randomly (on an invisible circle) in the image. There are also $2n$ edges between these nodes that form either one cycle of size $2n$ or two cycles of size $n$. The task is to determine whether the graph is connected (one cycle, label 1) or not (two cycles, label 0). See Figure 2 for an example. In this task, one has to reason over at least $n$ nodes and the connections between them to determine the label correctly, as any $n-1$ nodes provide no information on whether there are two cycles or one. Thus, one can simply increase the complexity of this task by increasing $n$.
    • Strings task. In order to further increase the visual complexity, we consider a dataset consisting of random strings. In each sample, there are either two closed strings or one longer closed string. The dataset generation process for these curves is similar to the cycles task above, except that in the strings we do not make the (anchor) nodes visible and also connect them using third-degree Bézier curves which produces continuous strings (see Figure 2). Similar to the cycles task, one can increase the complexity of this task by increasing the number of invisible anchor points $2n$, which leads to longer, more entangled strings.
  • Maze solvability. We also consider a maze task in which there are always two connected components, and we have a start/source point (shown in blue) and an end/sink point (shown in red). The source and sink are in the same connected component or not equiprobably. The task is to determine whether they are connected (label 1) or not (label 0). We provide this dataset in a rectangular and a circular version to increase the visual complexity. Examples can be seen in Figure 2. To adjust the complexity of maze datasets, one can modify the size of the maze and hence the number of cells, the size of the components, and the distance between the source and sink (if connected).
  • Image Pointer Value Retrieval (PVR). In our image pointer value retrieval (PVR) task, similar to Zhang et al. (2021b); Abbe et al. (2022a), each image is a grid of fixed size $n\times n$ (for $n\leq 10$) where there is an MNIST (LeCun et al., 2010) digit acting as a pointer in the top left cell of the grid pointing to one of the next $n-1$ rows. For a fixed $k$, in each of the next $n-1$ rows of the grid, we have $k$ images where each image is sampled uniformly from the plane and car categories of the CIFAR-10 (Krizhevsky et al., 2009) dataset. The label of an image is given by the parity of the number of occurrences of the leftmost object in the indicated row. [^2] For even $k$, the label is equal to the parity of the number of cars, which is the same as the parity of the number of planes. Note that the globality of this task is at least $k+1$, as one has to use the pointer and also all the $k$ images of the corresponding row to have non-zero information about the label (as any $k-1$ images of a row have no mutual information with the target since the parity function can flip depending on the single unseen image). Thus, one can also easily adjust the globality of the task by varying $k$. Figure 2 shows a PVR task with $n=7$ and $k=4$.
For each task, there exists a natural chain-of-sketch (a single frame or a sequence of frames) that uncovers the underlying reasoning behind the label. For the maze, similar to what humans do, we can start coloring from the source cell (i.e., the cell in blue) to see which areas are reachable until reaching the sink cell or the end of the maze region, similar to performing a breadth-first search (BFS). This coloring is similar to what humans would naturally do by following the paths from the starting point to see which one (if any) leads to the sink cell. For the cycles task, we can use a similar idea: we can start by coloring one node, and then color all of the nodes that are connected to this node (i.e., either half or all of the graph). Analogously, for the strings task, the CoS would be coloring one of the strings if there are two strings or coloring the whole string if there is only one. To disambiguate which cycle/string to color, we always color the cycle/string that passes through the rightmost (anchor) node. For the PVR task, we can keep only the row indicated by the pointer and then compute the parity from left to right by checking whether each object belongs to the same class as the leftmost object, and updating the parity accordingly. See Figure 2 for single-frame CoS examples for different tasks.
Note that the CoS can have a single frame format where the full sketch (e.g., with all the coloring done) is shown. The CoS can also be generated in multiple frames, i.e., consecutive frames that lead to the final sketch. This CoS is again analogous to what humans do as they progressively create the full sketch. For example, this could be coloring a distance of $10$ when doing the search for the maze problems; coloring (up to) two anchor nodes in the cycles and strings tasks; and checking one additional object in the PVR task. An example of doing so for the cycles task is depicted in Figure 3. See Appendix F for CoS frames for other datasets.
@@ 2 Global visual reasoning datasets
Figure 3: An example of the cycles task with a CoS containing several frames. The input image is presented on the left, followed by different frames of the CoS from left to right, ending with the complete sketch.
2.1 Globality in the visual domain
@@
Recently, Abbe et al. (2024) proposed the notion of globality degree to explain why tasks requiring global reasoning are hard for Transformers to learn, and to understand the effectiveness of scratchpads (Nye et al., 2021) and chain-of-thought (Wei et al., 2023) techniques in the text domain. For input tokens $X_{1},\ldots,X_{n}$ and output $Y$, the globality degree of a task is defined as the minimum number of tokens $k$ such that there exist $k$ tokens $X_{i_{1}},\ldots,X_{i_{k}}$ that along with the histogram of tokens $\hat{P}_{X}$ [^3] In text, histogram refers to reporting how many times each token is appearing regardless of its position (similar to the bag of words). provide significant information on the target $Y$, i.e., $I(X_{i_{1}},\ldots,X_{i_{k}},\hat{P}_{X};Y)=n^{-O_{n}(1)}$ where $I$ is the mutual information. It is further conjectured, with empirical support, that the learning complexity of tasks increases with their globality degree, and Transformers can only learn tasks with a constant globality degree efficiently (using a model of polynomial size and a polynomial number of iterations). We extend this definition to vision tasks learned with models like vision Transformers (ViTs) (Dosovitskiy et al., 2020), using a finer-grained formulation that integrates the amount of information.
@@
Globality degree (in the visual domain; see Appendix E). Assume images are partitioned into patches $X_{1},\ldots,X_{n}$. We define the globality degree with threshold $\alpha$ of a task as the minimum number $k$ such that there exist patches $X_{i_{1}},\ldots,X_{i_{k}}$ that satisfy $I(X_{i_{1}},\ldots,X_{i_{k}};Y)\geq\alpha$ where $I$ is the mutual information.
In words, this is the least number of patches $k^{*}$ required to obtain an $\alpha$-mutual information with the target. The higher $\alpha$, the more informative these patches are about the target, and the lower $k^{*}$ (for the same $\alpha$), the less global the task is. Given the arguments given by Abbe et al. (2024), we expect the learning complexity of the task to scale with $\frac{n^{k^{*}}}{\alpha}$ where $k^{*}$ is the globality degree. Therefore, the tasks become easier to learn as they have smaller globality degrees (e.g., $k^{*}=O_{n}(1)$) and harder to learn as their globality degrees increases. We further discuss the theoretical aspects in Appendix E.
Remark 1.
@@
Note that a target being “local” by having a low globality degree $k^{*}$ for a significant $\alpha$ does not mean that the target depends only on these few $k^{*}$ patches. The target can still depend on all the patches. What it means is that these $k^{*}$ patches are sufficient in order to obtain non-trivial information about the target, i.e., the model can achieve a non-trivial accuracy using these patches and thus weakly learn the task (this weak learning may as well provide a starting point for strong learning). We further clarify this using an experiment in Appendix D.1.
@@
According to the definition above, vision tasks such as classical image classification are “local” as a few patches often provide significant information on the class (e.g., having a patch containing a dog’s ear significantly increases the likelihood of predicting the dog class). Whereas in our proposed datasets, seeing a few patches often provides no information about the label. Hence, our proposed datasets have a high globality degree, or in short, are “global”. For example, in the maze examples, seeing just a few patches from the maze does not help determine the label. Similarly, for the cycles task of size $2n$, if one only sees the connections between $n-1$ nodes, one has no information about the label.
@@
Empirical validation. To further support our claim that mainstream vision tasks are local while our proposed tasks are not, we conduct the following experiment. For each sample in a given dataset, we mask the patches with probability $p$ at both training and inference and measure the performance of the model for different values of $p$. We perform this experiment on the cycles 12 task and ImageNet. Since the cycles 12 task is not learnable from scratch (see Figure 5), we start from a CLIP (Radford et al., 2021) pre-trained ViT-L/14 checkpoint (Fang et al., 2023) for both. The results are shown in Figure 5. We observe that the model demonstrates good performance on the ImageNet dataset even when 90% of the image is masked while it cannot learn the cycles 12 task once 30% (or more) of the image is masked. [^4] We use min-max normalized accuracy in the plot, including the random baseline for normalization. The latter shows that the cycles 12 task is a high globality dataset where one needs on average at least 70% of the patches to gain minimal information on the label while ImageNet is a local task where weak learning is possible with only 10% of the patches.
@@
Figure 4: Experimental evidence that cycles 12 is more global compared to common computer vision benchmarks (e.g. ImageNet). Cycles 12 quickly becomes hard to learn when more patches are masked, while ImageNet remains well above random accuracy.